OpenStudy (anonymous):

Given tanB = 15/9 find angle B in degrees. Round your answer to the nearest hundredth. Explain/ show work pleasee

OpenStudy (anonymous):

Are you allowed to use a calculator?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

If so, then note that \(\tan B = \dfrac{15}{9} \implies \tan^{-1}(\tan B)=\tan^{-1}\left(\dfrac{15}{9}\right) \implies B = \tan^{-1}\left(\dfrac{15}{9}\right)\). Most calculators should have the capability to compute this value. What kind of calculator are you using?

OpenStudy (anonymous):

google lol

OpenStudy (anonymous):

lol, that's fine. Just type in "inverse tangent of (15/9) in degrees" into google and it will give you an approximate value.

OpenStudy (anonymous):

59.036243?

OpenStudy (anonymous):

Yes, so what would that be to the nearest hundredth?

OpenStudy (anonymous):

59.04 thank you(:

OpenStudy (anonymous):

You're welcome! :)